A charge 4.99 nC is placed at the origin of an xy-coordinate system, and a charge -1.98 nC is placed on the positive x-axis at x = 3.96 cm . A third particle, of charge 6.04 nC is now placed at the point x-3.96 cm , y-3.05 cm

1.) Find the x-component and y-component of the total force exerted on the third charge by the other two.

2.) What is the magnitude of the total force acting on the third charge.

3.) What is the direction of the total force acting on the third charge.
Answer #1

Fя в D = 3-962+3.052 = 24.984! x = 24.9841X154 mL 3.+50= y F- -- - -> Fi Coso 82 8 = 3.05cm 23.05xlom o= tan (305) 0 = 37.60